OpenStudy (kendricklamar2014):

Find the simplified form of the expression. Give your answer in Scientific Notation. (8 * 10^4)(9 * 10^-8) A. 7.2 * 10^-5 B. 7.2 * 10^-3 C. 72 * 10^-4 D. 7.2 * 10^-31

OpenStudy (kendricklamar2014):

@Michele_Laino

OpenStudy (kendricklamar2014):

8 * 9 = 72

OpenStudy (michele_laino):

ok!

OpenStudy (kendricklamar2014):

10^4 * 10^-8 = 10^-4

OpenStudy (michele_laino):

ok!

OpenStudy (kendricklamar2014):

So with that ^ it comes out to be: 72 * 10^-4

OpenStudy (michele_laino):

yes it is correct, nevertheless it is not in scientific notation

OpenStudy (michele_laino):

please keep in mind that 72= 7.2 * 10^1

OpenStudy (kendricklamar2014):

Yes, would it be B

OpenStudy (michele_laino):

that's right!

OpenStudy (kendricklamar2014):

Thanks @Michele_Laino

OpenStudy (michele_laino):

Thanks! :)